# ProveRAG: Provenance-Driven Vulnerability Analysis with Automated Retrieval-Augmented LLMs

### v2

@@ Abstract
In cybersecurity, security analysts face the challenge of mitigating newly discovered vulnerabilities in real-time, with over 300,000 Common Vulnerabilities and Exposures (CVEs) identified since 1999. The sheer volume of known vulnerabilities complicates the detection of patterns for unknown threats. While LLMs can assist, they often hallucinate and lack alignment with recent threats. Over 25,000 vulnerabilities have been identified so far in 2024, which are introduced after popular LLMs’ (e.g., GPT-4) training data cutoff. This raises a major challenge of leveraging LLMs in cybersecurity, where accuracy and up-to-date information are paramount. In this work, we aim to improve the adaptation of LLMs in vulnerability analysis by mimicking how analysts perform such tasks. We propose ProveRAG, an LLM-powered system designed to assist in rapidly analyzing CVEs with automated retrieval augmentation of web data while self-evaluating its responses with verifiable evidence. ProveRAG incorporates a self-critique mechanism to help alleviate omission and hallucination common in the output of LLMs applied in cybersecurity applications. The system cross-references data from verifiable sources (NVD and CWE), giving analysts confidence in the actionable insights provided. Our results indicate that ProveRAG excels in delivering verifiable evidence to the user with over 99% and 97% accuracy in exploitation and mitigation strategies, respectively. This system outperforms direct prompting and chunking retrieval in vulnerability analysis by overcoming temporal and context-window limitations. ProveRAG guides analysts to secure their systems more effectively while documenting the process for future audits.
1 Introduction
@@
In today’s fast-paced cybersecurity landscape, security analysts are continuously challenged to understand and mitigate newly discovered threats in real-time [24]. Common Vulnerabilities and Exposures (CVEs) [22] serve as a critical framework for identifying and detecting known vulnerabilities. With over 300,000 CVEs cataloged since 1999, the sheer volume makes it increasingly difficult and time-consuming for analysts to detect patterns and respond effectively to new and emerging threats, particularly when faced with unknown vulnerabilities. Public vulnerability databases, such as the NIST National Vulnerability Database (NVD) [23] and privately maintained repositories, predominantly rely on labor-intensive manual processes to analyze and categorize vulnerabilities. This reliance on human effort can lead to delays in identifying and addressing potential threats, making it difficult to keep pace with the rapidly evolving threat landscape, where timely analysis is critical for effective mitigation [24]. For example, the CVE-2024-0302 is a critical vulnerability that, if left unchecked, could lead to deserialization of untrusted data. To mitigate this vulnerability, fields should be made transient to be protected from deserialization. However, protecting an enterprise network necessitates discovering which vulnerabilities are relevant by combing through over 300,000 threats aggregated in these databases with more added each day. Separate mitigation databases, such as Common Weakness Enumeration (CWE) [21], must be explored to identify, implement, and deploy mitigations to the system. Due to the sheer volume of information, manually identifying and mitigating threats is challenging. In addition, automating the identification and mitigation process is essential for reducing the risk and enhancing overall resilience.
@@
Large Language Models (LLMs) have been widely adopted in cybersecurity applications [20, 33, 3, 4, 12]. They have been particularly effective in assisting with the analysis of Cyber Threat Intelligence [20, 26], enhancing log-based anomaly detection [12, 27], supporting LLM-assisted attacks [31, 7, 4], and improving vulnerability detection [3, 33]. However, prior work has identified three aspects that limit the broader adoption of LLMs to cybersecurity applications [8, 28, 37, 15], which we aim to address in this work. These limitations are: 1) Hallucination: the propensity to generate misleading information that appears plausible but is factually incorrect. Some works have shown the consequences of this problem [8, 28]. 2) Temporal Knowledge: the failure of LLMs to respond accurately when they are exposed to data that comes after their training data’s cutoff [38, 37]. In cybersecurity, daily emerging vulnerabilities—over 25,000 identified in 2024 alone—pose serious risks, especially since they appeared after LLMs’ training data cutoff. Retrieval Augmented Generation (RAG) techniques have been proposed to address this problem [2]. However, the adoption of RAG for out-of-distribution data often leads to 3) Limited Context Windows: which restrict the number of tokens that can be provided to prompt the model, thereby limiting RAG-based temporal alignment approaches [15]. These limitations constitute a major challenge of leveraging LLMs in cybersecurity, where accuracy, time, and up-to-date information is essential.
@@
In this work, we aim to improve the adaptation of LLMs in threat analysis. To do so, we propose ProveRAG, an advanced LLM-powered system designed to assist security teams in rapidly analyzing CVEs with verifiable evidence. ProveRAG goes beyond merely identifying vulnerabilities; it generates comprehensive exploitation steps and provides actionable mitigation strategies. What sets ProveRAG apart is its ability to cross-reference, rationalize, and validate information from authoritative sources like the NVD and CWE, ensuring that the insights it delivers are accurate and trustworthy. By doing so, we demonstrate that ProveRAG exhibits fewer instances of omission and hallucination through self-critiquing its own response. A key innovation in ProveRAG is its incorporation of provenance into its analysis. By documenting the source of every piece of information, ProveRAG mitigates the risk of inaccuracies or hallucinations common in LLMs and gives security analysts the confidence to make swift, informed decisions.
@@
We test ProveRAG on 2024 CVEs, and we compare with two different baseline approaches: 1) prompt-only and 2) chunking retrieval during generation, and we compare it to our proposed summarizing technique (integrated into ProveRAG). The prompt-only technique is directly asking the LLM details about a specific CVE exploitation and mitigation strategies with data after its training cutoff knowledge. For the chunking technique, we enhance information retrieval by splitting and embedding the content into smaller chunks. The top-k results are retrieved and fed as context into the LLM’s prompt, which allows the model to answer the user’s query with the most relevant content. However, this approach requires a very large context window and can include redundant information that might mislead the LLM.
@@
Therefore, we design a summarizing technique to process CVEs, inspired by [14, 10]. We start by exploring the NVD website and summarize the content with respect to a specific CVE-ID. Next, we move to the associated CWE link, and further to the relevant references (hyperlinks) to provide summaries with respect to the exploitation and mitigation questions (emulating the typical workflow of a security analyst). A key advantage of the summarizing technique is its efficiency in handling large-scale data without relying extensively on document chunking, which enhances both retrieval quality and the accuracy of responses (while handling the context window limitation). This makes the system particularly suited for addressing more complex cybersecurity queries — such as those related to CVE mitigation that require integrating more external information.
There is no easy way to determine whether a generated response is accurate or trustworthy. ProveRAG has the ability to provide verifiable evidences (TPs), detect hallucinations (FPs) and failed retrievals (FNs) for the LLM responses. This is called provenance where we carefully designed a structured, explicit, concise prompt with the chain-of-thought technique [35] to push the LLM into providing evidence for its responses. Note that we ensure that the LLM for ProveRAG (gpt-4o-mini [25]) do not miss any essential information when finding evidence by utilizing almost the entire context window (128,000 tokens) for provenance.
@@
Moreover, we demonstrate that by providing the LLM with highly-structured context (Aqua Vulnerability Database [29]), ProveRAG’s performance increases significantly. Our testing demonstrates that ProveRAG achieves outstanding detection accuracy, with 99% for exploitation information and 97% for mitigation strategies. ProveRAG not only enables more reliable decision-making (compared to the chunking technique with over 30% increase in accuracy in providing mitigation strategies) but also meticulously documents its processes, ensuring a robust audit trail for future reviews. For evaluation, we ensure comprehensive coverage of all relevant information for each CVE by incorporating external resources beyond the NVD and CWE. Finally, we show that the proposed ProveRAG framework supports generalizability across different models by comparing OpenAI’s gpt-4o-mini and Meta’s Llama-3.1-8B [17]. This makes ProveRAG an invaluable system for security analysts, empowering them to safeguard their systems with greater efficiency and accuracy. The contributions of our work is as follows:
@@
  • We propose ProveRAG, an LLM-powered system that generates information from external sources with a summarizing technique, while self-criticising its responses with rationale by providing evidence from source URLs (i.e., provenance).
  • We demonstrate the effectiveness of ProveRAG for handling data that is out of the scope of LLMs’ knowledge cutoff with a very high detection rate (i.e., TPs).
  • We show that ProveRAG is able to show proof for its responses (TPs), detect hallucinations (FPs), as well as demonstrating failed retrievals (FNs), which alleviates the hallucination and omission limitations of state-of-the-art LLMs.
  • We demonstrate that the ProveRAG framework enhances vulnerability analysis by incorporating an additional LLM (Llama-3.1-8B), thereby supporting the generalizability of the proposed approach.
  • We show that ProveRAG relaxes the limitations imposed by context windows, which makes it suitable for answering more complex queries (e.g., mitigation-related) that require large-scale additional up-to-date information.
2 Background & Related Work
2.1 Large Language Models
@@
2.2 Retrieval Augmented Generation
@@
2.3 LLMs in Vulnerability Assessment
There are multiple works that integrated the use of LLMs in vulnerability assessment [13, 3, 5, 16]. Anton C. et al. [3] used GPT models to identify CWE vulnerabilities in Java code and revealed that LLMs struggle with detecting vulnerable code. In contrast, Avishree K. et al. [13] demonstrated with multiple prompt strategies that LLMs can outperform deep learning models in vulnerability detection when carefully prompted, which indicates the importance of prompt engineering. Building on this, Lu et al. [16] introduced GRACE, a framework using in-context learning and graph-structured information with LLMs, which showed improvements in vulnerability detection on C/C++ datasets. Du et al. [5] introduced Vul-RAG, LLM-based vulnerability detection technique that constructs a comprehensive knowledge base by extracting multi-dimensional insights from existing CVE instances. However, the reliance on human analysis in their approach makes it a manual time-consuming work and limits its efficiency in real-time vulnerability detection. In addition, PentestGPT [4] has shown capabilities in assisting in penetration testing. While it does not address the same problem this paper focuses on, the authors noted that LLMs are constrained by limited token sizes (i.e., context window), preventing long-term memory retention. This work builds on this observation and develop efficient RAG to bring LLM’s attention to relevant contents.
3 Motivation & Problem Formulation
@@
Traditional approaches to threat analysis typically involve manual assessment and the use of static databases, which can be both time-consuming and prone to human error. Moreover, while existing systems can assess vulnerabilities, they often fall short in providing detailed, actionable insights necessary for effective mitigation. This gap in the process leaves security teams with a critical need for more advanced systems capable of assessing threats, generating comprehensive mitigation scenarios, and documenting evidences of their recommendations. The motivation behind this work stems from these challenges. We seek to address the limitations by answering the following research questions:
@@
RQ1: How to address the temporal limitations of LLMs for continuously emerging cybersecurity threats? One of the key problems we address is the temporal limitation of LLMs. With the constant emergence of new vulnerabilities, it is crucial that security systems remain effective even when dealing with data that falls outside the LLM’s training window. Our system, ProveRAG, is designed to overcome this challenge by integrating up-to-date information from verified resources, which helps to assess vulnerabilities that have been discovered post-training.
@@
RQ2: How to manage information overload in the context of RAG-based LLMs for complex queries? Another significant problem we aim to solve is the issue of information overload (i.e., arising from context window limitations). This problem of inability of LLMs to process massive amounts of information and identify relevant information can lead to less effective responses, as discussed in literature [15]. While more information is often seen as beneficial, our research shows that in the context of LLM-powered vulnerability analysis, an abundance of data can lead to inefficiencies (when chunking technique is used). The summarization technique we encountered in ProveRAG, can alleviate this issue. This technique can make the system especially effective in handling more complex cybersecurity issues, such as CVE mitigation queries that require more attention on additional relevant resources.
@@
RQ3: How to enhance the accuracy and trustworthiness of threat analysis while addressing issues related to hallucination and omission errors? One other key motivation for developing ProveRAG is the critical requirement to enhance the accuracy and trustworthiness of threat analysis by mitigating issues related to hallucination and omission errors. ProveRAG addresses this by integrating authoritative sources such as NVD and CWE into its framework. This integration ensures that the recommendations provided by ProveRAG are grounded in reliable and verifiable data, which we call provenance. To leverage the capabilities of LLMs, ProveRAG utilizes a structured, explicit, and concise prompt designed with a chain-of-thought technique [35]. This approach compels the LLM to substantiate its responses with evidence, enhancing the credibility of the output. The system’s commitment to provenance is highlighted by its meticulous documentation of the sources of information it uses, creating a transparent and auditable trail. Based on our evaluation, this feature not only bolsters the analytical accuracy of ProveRAG but also fosters trust among security analysts by enabling them to verify the system’s recommendations.
4 ProveRAG Methodology

### v3

@@ Abstract
In cybersecurity, security analysts constantly face the challenge of mitigating newly discovered vulnerabilities in real-time, with over 300,000 vulnerabilities identified since 1999. The sheer volume of known vulnerabilities complicates the detection of patterns for unknown threats. While LLMs can assist, they often hallucinate and lack alignment with recent threats. Over 40,000 vulnerabilities have been identified in 2024 alone, which are introduced after most popular LLMs’ (e.g., GPT-5) training data cutoff. This raises a major challenge of leveraging LLMs in cybersecurity, where accuracy and up-to-date information are paramount. Therefore, we aim to improve the adaptation of LLMs in vulnerability analysis by mimicking how an analyst performs such tasks. We propose ProveRAG, an LLM-powered system designed to assist in rapidly analyzing vulnerabilities with automated retrieval augmentation of web data while self-evaluating its responses with verifiable evidence. ProveRAG incorporates a self-critique mechanism to help alleviate the omission and hallucination common in the output of LLMs applied in cybersecurity applications. The system cross-references data from verifiable sources (NVD and CWE), giving analysts confidence in the actionable insights provided. Our results indicate that ProveRAG excels in delivering verifiable evidence to the user with over 99% and 97% accuracy in exploitation and mitigation strategies, respectively. ProveRAG guides analysts to secure their systems more effectively by overcoming temporal and context-window limitations while also documenting the process for future audits.
Index Terms:
ProveRAG, LLM, Provenance, CVE, CWE, RAG, Vulnerability, Self-Critique
[…]
=-21pt
I Introduction
@@
In today’s fast-paced cybersecurity landscape, security analysts are continuously challenged to understand and mitigate newly discovered threats in real-time [1]. Common Vulnerabilities and Exposures (CVEs) [2] serve as a critical framework for identifying and detecting known vulnerabilities. With over 300,000 CVEs cataloged since 1999, the sheer volume makes it increasingly difficult and time-consuming for analysts to detect patterns and respond effectively to emerging threats. Public vulnerability databases, such as the NIST National Vulnerability Database (NVD) [3], predominantly rely on labor-intensive manual processes to analyze and categorize vulnerabilities. This reliance on human effort can cause delays in threat identification and response, which makes it difficult to keep pace with the rapidly evolving threat landscape [1].
For example, the CVE-2024-0302 is a critical vulnerability that, if left unchecked, could lead to deserialization of untrusted data. To mitigate this vulnerability, fields should be made transient to be protected from deserialization. However, protecting an enterprise network necessitates discovering which vulnerabilities are relevant by combing through over 300,000 threats aggregated in these databases, with more added each day. Separate mitigation databases, such as the Common Weakness Enumeration (CWE) [4], must also be explored to identify, implement, and deploy mitigations to the system. Due to the sheer volume of information, manually identifying and mitigating threats is challenging, and automating the identification and mitigation process is essential to reduce risk and improve overall resilience.
Large Language Models (LLMs) have been widely adopted in cybersecurity applications [5, 6, 7, 8]. They have been particularly effective in assisting with the analysis of Cyber Threat Intelligence [5, 9], supporting LLM-assisted attacks [10, 11, 8], and improving vulnerability detection [7, 6].
@@
However, prior work has identified three aspects that limit the broader adoption of LLMs to cybersecurity applications [12, 13, 14, 15], which we aim to address in this work. These limitations are: 1) Hallucination: the tendency to generate misleading information that appears plausible but is factually incorrect. Some works have shown the consequences of this problem [12, 13]. 2) Temporal Knowledge: the failure of LLMs to respond accurately when they are exposed to data that come after the cutoff point of their training data [16, 14]. In cybersecurity, daily emerging vulnerabilities—over 40,000 identified in 2024 alone—pose serious risks, especially if they appear after LLMs’ training data cutoff. Retrieval Augmented Generation (RAG) techniques have been proposed to address this problem [17] by retrieving relevant external information. However, adoption of RAG often leads to 3) Limited Context Windows: which restrict the number of tokens that can be provided to prompt the model, thus limiting RAG-based temporal alignment approaches [15]. These limitations constitute a major challenge of leveraging LLMs in cybersecurity, where accuracy, time, and up-to-date information is essential.
@@
In this work, we aim to improve the adaptation of LLMs in vulnerability analysis. To do so, we propose ProveRAG, an advanced LLM-powered system designed to assist security teams in rapidly analyzing CVEs with verifiable evidence. What sets ProveRAG apart is its ability to cross-reference, rationalize, and validate information from authoritative sources, such as NVD and CWE. By doing so, we demonstrate that ProveRAG exhibits fewer instances of omission and hallucination through self-critiquing its own response.
A key innovation in ProveRAG is its incorporation of provenance into its analysis. By documenting the source of every piece of information, ProveRAG reveals the hallucinations and inaccuracies common in LLMs’ responses and gives security analysts the confidence to make informed decisions. For example, when suggesting mitigation strategies for vulnerabilities like CVE-2024-0302, ProveRAG reveals pieces of data from trusted databases (such as NVD and CWE) that led to its generated response. This provides a clear audit trail that can be reviewed and verified, thus fostering trust and enabling effective action.
@@
We test ProveRAG on 2024 CVEs, and we compare with two different baseline approaches: 1) prompt-only and 2) chunking retrieval during generation We compare these baselines to our proposed summarizing technique (integrated into ProveRAG). The prompt-only technique is directly asking the LLM details about a specific CVE exploitation and mitigation strategies with data after its training cutoff knowledge. For the chunking technique, we enhance information retrieval by splitting and embedding the retrieved content into smaller chunks. However, this approach requires a very large context window and can include redundant information that might mislead the LLM.
@@
Therefore, we design a summarizing technique to process CVEs, inspired by [18, 19] in an automatic way by mimicking how security analysts perform vulnerability analysis. To do so, we start by exploring the NVD website and summarize the content with respect to a specific CVE-ID. Next, we move to the associated CWE link, and further to the relevant references (hyperlinks) to provide summaries with respect to the exploitation and mitigation questions (emulating the typical workflow of a security analyst). A key advantage of the summarizing technique is its efficiency in handling large-scale data without relying extensively on document chunking, which enhances both retrieval quality and the accuracy of responses (while handling the context window limitation). This makes the system particularly suited to better guide analysts in a timely manner while addressing more complex cybersecurity queries - such as those related to CVE mitigation that require integrating more external information.
ProveRAG has the ability to provide verifiable evidence (TPs), while also revealing hallucinations (FPs) and failed retrievals (FNs) to determine whether a generated response is accurate and trustworthy (the provenance part of the proposed framework). Furthermore, we demonstrate that by providing the LLM with highly-structured context (Aqua Vulnerability Database [20]), ProveRAG’s performance increases significantly. Our testing shows that ProveRAG achieves outstanding detection accuracy, with 99% for exploitation and 97% for mitigation strategies.
@@
ProveRAG not only enables more reliable decision-making but also documents its processes to ensure a robust audit trail for future reviews. For evaluation, we ensure comprehensive coverage of all relevant information for each CVE by incorporating external resources beyond the NVD and CWE. Finally, we show that the proposed ProveRAG framework generalizes across different models by comparing OpenAI’s gpt-4o-mini and Meta’s Llama-3.1-8B [21]. This allows security analysts to rapidly navigate the massive amounts of data to identify relevant cybersecurity vulnerabilities applicable to their enterprise system and develop appropriate mitigation strategies. The contributions of our work are as follows:
@@
  • We propose ProveRAG, an LLM-powered system developed to self-critique its generated responses with rationale, by providing pieces of evidence from verifiable sources that led to its generation (i.e., provenance).
  • We apply ProveRAG to perform vulnerability analysis, a critical application that requires retrieval and verification of new, unseen knowledge.
  • We demonstrate that ProveRAG not only can show proof for its responses (TPs), but also reveals hallucinations (FPs) and failed retrievals (FNs) with evidence.
  • We show that ProveRAG effectively respond to unseen data while relaxing the limitations imposed by context windows, leading to high #TPs (99% accuracy) and low #FPs and #FNs, even when answering complex queries (e.g., vulnerability mitigation).
  • We evaluate ProveRAG on two LLMs, gpt-4o-mini and Llama-3.1-8B, to support the generalizability of the proposed framework.
II Background & Related Work
II-A Large Language Models
@@
II-B Retrieval Augmented Generation
@@
II-C LLMs in Vulnerability Analysis
There are multiple works that integrated the use of LLMs in vulnerability analysis [31, 7, 32, 33]. Anton C. et al. [7] used GPT models to identify CWE vulnerabilities in Java code and revealed that LLMs struggle with detecting vulnerable code. In contrast, Avishree K. et al. [31] demonstrated with multiple prompt strategies that LLMs can outperform deep learning models in vulnerability detection when carefully prompted, which indicates the importance of prompt engineering. Building on this, Lu et al. [33] introduced GRACE, a framework using in-context learning and graph-structured information with LLMs, which showed improvements in vulnerability detection on C/C++ datasets. Du et al. [32] introduced Vul-RAG, LLM-based vulnerability detection technique that constructs a comprehensive knowledge base by extracting multi-dimensional insights from existing CVE instances. However, the reliance on human analysis in their approach makes it a manual time-consuming work and limits its efficiency in real-time vulnerability detection. In addition, PentestGPT [8] has shown capabilities in assisting in penetration testing. While it does not address the same problem this paper focuses on, the authors noted that LLMs are constrained by limited token sizes (i.e., context window), preventing long-term memory retention. This work builds on this observation and develop efficient RAG to bring LLM’s attention to relevant contents.
III Motivation & Problem Formulation
@@
Vulnerability analysis involves evaluating a system’s configuration to identify potential security risks. By assessing attack surfaces and risk factors, it helps uncover weak points in the system that require effective mitigation to prevent exploitation. Traditional approaches to vulnerability analysis typically involve manual assessment and the use of static databases, which can be both time-consuming and prone to human error. Moreover, while existing systems can assess vulnerabilities, they often fall short in providing detailed, actionable insights necessary for effective mitigation. This gap in the process leaves security teams with a critical need for more advanced systems capable of assessing threats, generating comprehensive mitigation scenarios, and documenting evidences of their recommendations. The motivation behind this work stems from these challenges. We seek to address these limitations by answering the following research questions:
@@
RQ1: How can the knowledge cutoff of LLMs be addressed in the context of continuously emerging cybersecurity threats? One of the key problems we address is the knowledge cutoff (temporal limitation) of LLMs. With the constant emergence of new vulnerabilities, it is crucial that security systems remain effective even when dealing with data that falls outside the LLM’s training window. Our system, ProveRAG, is designed to overcome this challenge by integrating up-to-date information from verified resources to effectively assess vulnerabilities discovered post-training. It does this by systematically cross-referencing to standardized security databases (e.g., NIST NVD) in a sequential manner and gathering relevant information, particularly for complex queries that require more in-depth and thorough retrieval, such as those related to mitigation strategies.
@@
RQ2: How to manage information overload in the context of RAG-based LLMs for complex queries? Another significant problem we aim to solve is the issue of information overload (i.e., arising from context window limitations). This problem of inability of LLMs to process massive amounts of information and identify relevant information can lead to less effective responses, as discussed in literature [15]. While more information is often seen as beneficial, our research shows that in the context of LLM-powered vulnerability analysis, an abundance of data can lead to inefficiencies (when chunking technique is used). The summarization technique we encountered in ProveRAG, can alleviate this issue. This technique can make the system effective in handling more complex cybersecurity issues, such as CVE mitigation queries that require more attention on additional relevant resources.
@@
RQ3: How to enhance the accuracy and trustworthiness of threat analysis while addressing issues related to hallucination and omission errors? One other key motivation for developing ProveRAG is the critical requirement to enhance the accuracy and trustworthiness of vulnerability analysis by mitigating issues related to hallucination and omission errors. ProveRAG addresses this by integrating authoritative sources such as NVD and CWE into its framework. This integration ensures that the recommendations provided by ProveRAG are grounded in reliable and verifiable data, which we call provenance. To leverage the capabilities of LLMs, ProveRAG utilizes an explicit and structured prompt designed with the chain-of-thought technique [34]. Based on our evaluation of ProveRAG (Sec. V-B2), this causes the LLM to substantiate all responses with direct evidence, enhancing the credibility of the output.
IV ProveRAG Methodology
The proposed ProveRAG methodology consists of two stages: The Generation and the Evaluation. An overview of each stage is in Figure 1. Note that the Retr. LLM, Gen. LLM, and Eval. LLM are the same model. For the remainder of this section, we discuss each stage in turn. First, in the generation module, the Retr. LLM will load web content from external sources about a CVE vulnerability—starting with NVD, followed by CWE, and then additional references (replication of a typical security analyst workflow in practice) —and summarizes the content with respect to exploitation and mitigation information. Next, the Gen. LLM will look at the summaries and generates detailed exploitation and mitigation strategies for that specific CVE. Furthermore, in the evaluation module, the Eval. LLM will critique its own response by looking at the content from the official and external (NVD, CWE, and hyperlinks in NVD) websites to show evidence on whether its response were accurate, hallucinated, or omitted critical information. Finally, the LLM will provide provenance by showing similar pieces of information from both the response and the retrieved context that led to its generation, followed by a rationale to ensure reliability and enhance decision-making. It is worth noting that we used the OpenAI’s gpt-4o-mini [23] model as the LLM for ProveRAG. This model’s training data knowledge is up to Oct. 2023, with 128,000 tokens for context window [23]. However, to show the proposed ProveRAG framework can be generalized across other LLMs, we compare the performance of gpt-4o-mini with Llama-3.1-8B [21] in Sec. V-E. The Llama-3.1-8B model is selected as it has the same context window of 128,000 tokens, and a knowledge cutoff date of Dec. 2023 (i.e., no 2024 CVEs are seen in training). Code and data are available at: https://github.com/RezzFayyazi/ProveRAG.
IV-A System Design
IV-A1 Generation Module:
{mybox}
[frametitle=Generation Prompt]
You are a cybersecurity expert. Consider the Relevant Information provided below and answer the Query.
Relevant Information: [content of NVD/CWE/Relevant Refs.]
Query: [CVE-xxx-xxx]
Given the specified CVE-ID, please provide detailed answers to the following questions:
1. How can an attacker exploit this vulnerability? Provide a step-by-step description.
2. What are the recommended mitigation strategies for this vulnerability?
The Relevant Information is gathered from the summarizing retrieval technique (to answer RQ1), and we will show in Sec. V-B1, that it is the better retrieval technique in vulnerability analysis. The Gen. LLM will look at the retrieved information for a particular CVE and will provide the response. In the prompt-only experiment, the Relevant Information part is removed as we directly query the LLM about a specific CVE. Note that we put the temperature hyperparameter (which assigns the degree of randomness in LLM’s output) to zero as we aim to get the most deterministic responses from the LLM.